please help
Write the equation of the parabola in vertex form.vertex (3, 3), point (–3, –33)

Answer by Edwin McCravy(20060)

please help
Write the equation of the parabola in vertex form.vertex (3, 3), point (–3, –33)

y = a(x - h)² + k

Substitute (h,k) = (3,3) and (x,y) = (-3,-33)

-33 = a(-3 - 3)² + 3

-33 = a(-6)² + 3

-33 = a(36) + 3

-33 = 36a + 3

-36 = 36a

-1 = a

So go back to the equation 

y = a(x - h)² + k

substitute -1 for a, 3 for h and 3 for k, and just leave x
and y as variables this time, and you have the equation:


y = (-1)(x - 3)² + 3

or just

y = -(x - 3)² + 3

Edwin